A 2.95 g sample of a compound consisting of carbon, hydrogen, oxygen, nitrogen, and sulfur was combusted in excess oxygen. This produced 1.66 g CO2 and 1.02 g H2O . A second sample of this compound with a mass of 4.55 g produced 4.66 g SO3 . A third sample of this compound with a mass of 9.00 g produced 3.63 g HNO3 . Determine the empirical formula of the compound. Enter the correct subscripts on the given chemical formula.
1
Expert's answer
2020-07-16T06:59:34-0400

moles of C = 1.66/44= 0.037

Moles of H = 1.02/18 = 0.05*2 = 0.113

Moles of O = 2.4/32 = 0.075

Moles of S = 4.66/64 = 0.0728

Moles of N = 3.63/63 = 0.057

Divide by the smallest to get the ratio

0.037/0.37: 0.113/0.037: 0.075/0.037: 0.0728/0.037: 0.057/0.037

1:3:2:2:2

empirical: CH3O2S2N2